# Weyl-Einsteinian-Cubic Gravity

### v1

@@ Appendix B The Pure Curvature Terms $\widehat{\Omega}_{(i)}$’s
$+16\alpha_{5}(R_{\mu\sigma}R^{\sigma\alpha\beta}{{}_{\nu}}-R^{\alpha\sigma}R_{\mu\sigma\nu}{{}^{\beta}})+4\alpha_{6}(R_{\mu\nu}R^{\alpha\beta}-R_{\mu}{{}^{\alpha}}R_{\nu}{{}^{\beta}}).$ (41)
@@
$\widehat{\Omega}^{(7)}_{\mu\nu}$ =$=\{8\alpha_{5}\square R_{\mu}{{}^{\alpha\beta\rho}}+8(-9\alpha_{2}+2\alpha_{5}-\alpha_{6})\nabla^{\rho}\nabla_{\mu}R^{\alpha\beta}+8(3\alpha_{3}-2\alpha_{5})\nabla^{\alpha}\nabla^{\beta}R^{\rho}{{}_{\nu}}+$
@@ C.3 Coefficient Functions of $\mathscr{O}^{\alpha}_{(3)\mu}$-Part:
wherein the explicit form of terms $\Xi^{(5)\alpha}_{\mu}$ and $\Xi^{(7)}_{\mu\sigma}$ are provided in (10). Also, the coupling cosntants $\kappa_{m(1)}$ (where $m=1\cdots 3$) and $\kappa_{1(3)}$ are consecutively given in (D.1) and (D.2).
@@ C.7 Coefficient Functions of $\mathscr{O}^{(7)}_{\mu\nu}$-Part:
($\triangle^{\alpha}_{(1)}A^{\beta}=\nabla^{\alpha}A^{\beta}-2A^{\alpha}A^{\beta},\,\,\,\,\,\,\,\triangle^{\alpha}_{(2)}A^{\beta}=\nabla^{\alpha}A^{\beta}+A^{\alpha}A^{\beta}$ (54)
^$\widehat{\boxdot}^{\mu\nu}A^{\alpha}=\nabla^{\mu}\nabla^{\nu}A^{\alpha}-2A^{\alpha}(\nabla^{\mu}A^{\nu})-2A^{\nu}(\nabla^{\mu}A^{\alpha})+A_{\lambda}R^{\mu\alpha\nu\lambda},$
where ^$\widehat{\boxdot}A_{\mu}=\widehat{\boxdot}_{\alpha}{{}^{\alpha}}A_{\mu}$.
@@ D.4 $\kappa_{j(6)}$ of $\Theta_{k}^{[6]j}A_{l}$
$\zeta_{17},\quad\kappa_{1(6)}=\kappa_{4(1)}+4\zeta_{23},\quad\kappa_{29(6)}=2\zeta_{6}+2\zeta_{12}-\zeta_{13}+72\alpha_{2}.$ (59)
@@ Appendix E Coefficient Functionals $\mathscr{X}_{[A]i}$’s of Gauge Field Equation
The explicit forms of coefficient functionals $\mathscr{X}_{[A]i}$ (with $i=1\cdots 11$) belonging to the gauge field equation ($\delta A^{\mu}$) in (22) are described successively as follows

### v2

@@ Appendix B The Pure Curvature Terms $\widehat{\Omega}_{(i)}$’s
[…]
@@
$\widehat{\Omega}^{(7)}_{\mu\nu}$ $=\{8\alpha_{5}\square R_{\mu}{{}^{\alpha\beta\rho}}+8(-9\alpha_{2}+2\alpha_{5}-\alpha_{6})\nabla^{\rho}\nabla_{\mu}R^{\alpha\beta}+8(3\alpha_{3}-2\alpha_{5})\nabla^{\alpha}\nabla^{\beta}R^{\rho}{{}_{\nu}}$+
@@ C.3 Coefficient Functions of $\mathscr{O}^{\alpha}_{(3)\mu}$-Part:
wherein the explicit form of terms $\Xi^{(5)\alpha}_{\mu}$ and $\Xi^{(7)}_{\mu\sigma}$ are provided in (12). Also, the coupling constants $\kappa_{m(1)}$ (where $m=1\cdots 3$) and $\kappa_{1(3)}$ are consecutively given in (D.1) and (D.2).
@@ C.7 Coefficient Functions of $\mathscr{O}^{(7)}_{\mu\nu}-$Part:
$\triangle^{\alpha}_{(1)}A^{\beta}=\nabla^{\alpha}A^{\beta}-2A^{\alpha}A^{\beta},\,\,\,\,\,\,\,\triangle^{\alpha}_{(2)}A^{\beta}=\nabla^{\alpha}A^{\beta}+A^{\alpha}A^{\beta}$ (56)
$\widehat{\boxdot}^{\mu\nu}A^{\alpha}=\nabla^{\mu}\nabla^{\nu}A^{\alpha}-2A^{\alpha}(\nabla^{\mu}A^{\nu})-2A^{\nu}(\nabla^{\mu}A^{\alpha})+A_{\lambda}R^{\mu\alpha\nu\lambda},$
where $\widehat{\boxdot}A_{\mu}=\widehat{\boxdot}_{\alpha}{{}^{\alpha}}A_{\mu}$.
@@ D.4 $\kappa_{j(6)}$ of $\Theta_{k}^{[6]j}A_{l}$
$\zeta_{17},\quad\kappa_{28(6)}=\kappa_{4(1)}+4\zeta_{23},\quad\kappa_{29(6)}=2\zeta_{6}+2\zeta_{12}-\zeta_{13}+72\alpha_{2}.$ (61)
@@ Appendix E Coefficient Functionals $\mathscr{X}_{[A]i}$’s of Gauge Field Equation
The explicit forms of coefficient functionals $\mathscr{X}_{[A]i}$ (with $i=1\cdots 11$) belonging to the gauge field equation in (24) are described successively as follows